Ahmad is twice as old as Bobby. John is 7 years younger than Ahmad. If the sum of their ages is 38, how old are the three boys?

Answer by checkley79(3341)

A=2B
J=A-7 OR 2B-7
A+B+J=38
(2B)+B+(2B-7)=38
5B-7=38
5B=38=7
5B=45
B=45/5
B=9 AGE OF BOBBY.
A=2*9=18 AGE OF AHMAD.
J=18-7=11 AGE OF JOHN.
PROOF:
18+9+11=38
38=38